Find the total number of oxygen atoms in 6 × 106
formula units of the complex
[Co(H2O)6]Cl2.
1
Expert's answer
2020-07-29T06:34:00-0400

Solution.

N(O in 1 [Co(H2O)6]Cl2) = 6

"N(O \ in \ 6\times 10^6 \ [Co(H2O)6Cl2]) = 6 \times 10^6 \times 6 = 36 \times 10^6"

Answer:

"N(O \ in \ 6\times 10^6 \ [Co(H2O)6Cl2]) = 36 \times 10^6"